### 13 Feb 2024 v1

# Forecasts for Constraining Lorentz-violating Damping of Gravitational Waves from Compact Binary Inspirals

Abstract
Violation of Lorentz symmetry can result in two distinct effects in the propagation of the gravitational waves (GWs). One is a modified dispersion relation and another is a frequency-dependent damping of GWs. While the former has been extensively studied in the literature, in this paper we concentrate on the frequency-dependent damping effect that arises from several specific Lorentz-violating theories, such as spatial covariant gravities, Hořava-Lifshitz gravities, etc. This Lorentz-violating damping effect changes the damping rate of GWs at different frequencies and leads to an amplitude correction to the GW waveform of compact binary inspiral systems. With this modified waveform, we then use the Fisher information matrix to investigate the prospects of constraining the Lorentz-violating damping effect with GW observations. We consider both ground-based and space-based GW detectors, including the advanced LIGO, Einstein Telescope, Cosmic Explorer (CE), Taiji, TianQin, and LISA. Our results indicate that the ground-based detectors in general give tighter constraints than those from the space-based detectors. Among the considered three ground-based detectors, CE can give the tightest constraints on the Lorentz-violating damping effect, which improves the current constraint from LIGO-Virgo-KAGRA events by about 8 times.
I INTRODUCTION
Since the landmark discovery of the first gravitational wave (GW) event, GW150914, resulting from the coalescence of two massive black holes by the LIGO-Virgo collaboration in 2015 LIGOScientific:2016aoc , the field of GW astronomy has rapidly evolved. To date, approximately 90 events have been meticulously identified by the LIGO-Virgo-KAGRA (LVK) scientific collaborations LIGOScientific:2018mvr ; LIGOScientific:2019lzm ; LIGOScientific:2020aai ; LIGOScientific:2021djp . On May 24, 2023, the advanced LIGO (aLIGO) initiated the Observing Run 4 (O4) project. Following LVK, the forthcoming third-generation ground-based GW detectors, such as Einstein Telescope (ET) Branchesi:2023mws and Cosmic Explorer (CE) Evans:2021gyd , are presently in the design phase, with a specific emphasis on detecting high-redshift GW events ($z>10$). Concurrently, a new cohort of space-based detectors (Taiji Ruan:2018tsw ; Wu:2018clg ; Hu:2017mde , TianQin Liu:2020eko ; Wang:2019ryf ; TianQin:2015yph ; Luo:2020bls ; Milyukov:2020kyg , and LISA Robson:2018ifk ; LISACosmologyWorkingGroup:2022jok ) is designed to explore the low-frequency GW signals ($f\sim 10^{-4}$ Hz). We anticipate that these detectors will play a crucial role in the era of GW astronomy LIGOScientific:2016lio ; LIGOScientific:2018dkp ; LIGOScientific:2017zic ; LIGOScientific:2019fpa ; LIGOScientific:2020tif .
General relativity (GR) remains the preeminent theory for explaining gravitational phenomena. Yet, it faces challenges in accounting for enigmatic concepts such as dark matter and dark energy, and reconciling them with quantum mechanics, particularly in the contexts of singularities and the quantization of gravity. To address these issues, a plethora of experiments have been devised to rigorously test GR’s predictions. Regrettably, the majority of these experiments have been limited to investigating the weak-field regime Hoyle:2000cv ; Everitt:2011hp ; Sabulsky:2018jma . GWs, one of the fundamental predictions of GR, are produced in the tumultuous environments of strong gravitational fields and interact only weakly with matter, making them pristine messengers of the dynamics of space-time. The detection of GWs, especially those originating from the coalescences of compact binary systems LIGOScientific:2020iuh ; LIGOScientific:2017vwq ; LIGOScientific:2020ibl , has thus heralded a new era for testing the robustness of GR under extreme conditions. These observations offer a powerful tool for probing the strong-field regime of gravity, potentially unlocking answers to the persistent questions that challenge the current understanding within the framework of GR.
In the theoretical realm, various modified theories of gravity Cognola:2006eg ; Copeland:2006wr ; Frieman:2008sn have been proposed to address challenges within GR. A subset of these theories has garnered significant attention for deviating from a fundamental principle of GR — the Lorentz invariance. At high-energy levels, it is widely believed that this invariance will be broken when gravity is quantized. Various modified gravity theories have been proposed to explore the nature of Lorentz violation in gravity, including the Einstein-Æther theory Jacobson:2000xp ; Eling:2004dk ; Jacobson:2007veq ; Li:2007vz ; Battye:2017zvv ; Zhang:2023kzs ; Liu:2021yev ; Zhu:2019ura , Hořava-Lifshitz theories of quantum gravity Horava:2009uw ; Takahashi:2009wc ; Wang:2012fi ; Zhu:2013fja , the standard model extension Kostelecky:2016kfm ; Bailey:2006fd ; Mewes:2019dhj ; Shao:2020shv ; Wang:2021ctl ; Kostelecky:2017zob ; Niu:2022yhr , and spatial covariant gravities Gao:2020qxy ; Gao:2020yzr ; Gao:2019twq ; Gao:2019liu ; Joshi:2021azw .
The violation of Lorentz symmetry in gravity can introduce deviations from GR in the propagation of GWs. These deviations manifest in two distinct ways, influencing the propagation behavior of GWs in the cosmological background. First, with Lorentz violation, the conventional linear dispersion relation of GWs can be modified into a nonlinear one, which in turn changes the phase velocities of GWs at different frequencies. This effect can arise from a large number of Lorentz-violating theories. Second, Lorentz violation can introduce frequency-dependent friction into the propagation equation of GWs, resulting in varying damping rates for GWs of different frequencies during their propagation. This effect normally arises from those theories with mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance, for example, the Hořava-Lifshitz gravity Colombo:2014lta , the spatial covariant gravities Gao:2019liu ; Zhu:2022uoq , etc.
Testing Lorentz symmetry of gravitational interaction by using the observational data from GW events in LIGO-Virgo-KAGRA catalogs and future GW detectors has been carried out in a lot of works, see Refs. LIGOScientific:2019fpa ; LIGOScientific:2020tif ; Mirshekari:2011yq ; Will:1997bb ; Zhu:2022uoq ; Gong:2021jgg and references therein. In most of these works, the effects due to the nonlinear dispersion relation have been extensively considered. Recently, the constraint on the Lorentz-violating damping effects from GW events in LIGO-Virgo-KAGRA catalogs was first obtained Zhu:2023wci .
In this paper, we detail the Lorentz-violating damping effects in the propagation of GW in a cosmological background. Decomposing the GWs into left-hand and right-hand circular polarization modes, we observe that the Lorentz-violating damping effects manifest through explicit modifications in the GW amplitude. We derive corrections to the waveform of the compact binary inspiral system accordingly. With this modified waveform, we use the Fisher information matrix (FIM), which is widely used in cosmology and astrophysics Yu:2023ico ; Jin:2023tou ; Shao:2023agv ; Jin:2023sfc ; Zhang:2023gaz ; Wu:2022jkf ; Song:2022siz ; Jin:2022qnj ; Wu:2022dgy ; Wang:2022oou ; Wu:2021vfz ; Jin:2021pcv ; Zhang:2021yof ; Yu:2021nvx ; Zhao:2019gyk ; Zhang:2019loq ; Wang:2019tto ; Zhang:2019ple to investigate the prospects of constraining the Lorentz-violating damping effect with GW observations of compact binary systems. We consider both ground-based and space-based GW detectors, including aLIGO, CE, ET, Taiji, TianQin, and LISA. Our results indicate that the ground-based detectors in general give tighter constraints than those from the space-based detectors. Among the considered three ground-based detectors, CE can give the tightest constraints on the Lorentz-violating damping effect, which improves the current constraint from LIGO-Virgo-KAGRA events by about 8 times Zhu:2023wci .
Our paper is organized as follows. In Sec. II, We present a very brief introduction to the Lorentz-violating damping effect and calculate the modified waveform of GWs of compact binary inspiral systems with the effect. Sec. III summarizes the application of the FIM for constraining the modified waveform parameters of GWs. The main results of our analysis are discussed in Sec. IV. Finally, Sec. V provides a summary and further discussion of our work in this paper.
II MODIFIED WAVEFORM OF GWS WITH LORENTZ-VIOLATING DAMPING EFFECT
In this section, we present a brief introduction to the modified waveform of GWs with Lorentz-violating damping effect. As we mentioned, such Lorentz-violating damping effect can modify the amplitude damping rates of the two tensorial modes of GWs, which arise from several specific Lorentz-violating theories of gravity, for instance, the spatial covariant gravities Gao:2019liu ; Zhu:2022uoq and Hořava-Lifshitz gravity Colombo:2014lta .
Figure 1: The noise spectral density of the six detectors considered in this paper. Both ground-based and space-based detectors are included in the picture.
II.1 Propagating equation of GWs with Lorentz-violating damping effects
Let us first investigate the propagation of GWs with Lorentz-violating damping effect on a flat Friedmann-Robertson-Walker spacetime. Treating this spacetime as a background, GWs can be described by the tensor perturbations of the metric, where the metric is expressed in the form of
$ds^{2}=a^{2}(\tau)\Big{[}-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\Big{]},$ (2.1)
where $a(\tau)$ is the scale factor of the expanding Universe and $\tau$ represents the conformal time. One can transform the conformal time $\tau$ to the cosmic time $t$ by $dt=a(\tau)d\tau$. Throughout this paper, we set the present expansion factor $a_{0}=1$. $h_{ij}$ denote GWs, which are transverse and traceless, i.e.,
$\partial^{i}h_{ij}=0=h^{i}_{i}.$ (2.2)
For later convenience, let us expand $h_{ij}$ over spatial Fourier harmonics,
$h_{ij}(\tau,x^{i})=\sum_{A=R,L}\int\frac{d^{3}k}{(2\pi)^{3}}h_{A}(\tau,k^{i})e^{ik_{i}x^{i}}e_{ij}^{A}(k^{i}),$ (2.3)
where $e_{ij}^{A}$ is the circular polarization tensor and obeys the following rules
$\epsilon^{ijk}n_{i}e^{A}_{kl}=i\rho_{A}e_{l}^{jA}$ (2.4)
with $\rho_{\rm R}=1$ and $\rho_{\rm L}=-1$.
To study the Lorentz-violating damping effect on the propagation of GWs, let us first write the modified propagation equation of motions of the two GW modes in the following parametrized form Zhu:2023wci ; Zhao:2019xmm ,
$h_{A}^{\prime\prime}+(2+{\bar{\nu}}+\nu_{A})\mathcal{H}h^{\prime}_{A}+(1+\bar{\mu}+\mu_{A})k^{2}h_{A}=0,$ (2.5)
where a prime denotes the derivative concerning the conformal time $\tau$ and ${\cal H}=a^{\prime}/a$. The four parameters, ${\bar{\nu}}$, $\nu_{A}$, ${\bar{\mu}}$, and $\mu_{A}$ label the new effects on the propagation of GWs arising from theories beyond GR. As mentioned in Ref. Zhu:2023wci , such parametrization provides a general framework for exploring possible modified GW propagations arising from a large number of modified theories of gravity. Different parameters correspond to different effects on the propagation of GWs. The parameters $\nu_{A}$ and $\mu_{A}$ represent the effects of parity violations, while the parameters ${\bar{\nu}}$ and ${\bar{\mu}}$, if frequency-dependent, can originate from other potential modifications involving Lorentz violations. For ${\bar{\nu}}$ and ${\bar{\mu}}$, the former provides an amplitude modulation of the GW waveform, while the latter one ${\bar{\mu}}$ determines the phase velocities of the GWs. In this paper, we will only concentrate on the case of the Lorentz-violating damping effect, and for this case one has
${\bar{\nu}}\neq 0,\;\;\nu_{A}=0,\;\;{\bar{\mu}}=0=\mu_{A}.$ (2.6)
Thus the propagation equation now can be written in the form of
$h^{\prime\prime}_{A}+(2+{\bar{\nu}})\mathcal{H}h^{\prime}_{A}+k^{2}h_{A}=0.$ (2.7)
With the Lorentz-violating damping effect, the only nonzero parameter ${\bar{\nu}}$ is generally frequency dependent. Thus, following Ref. Zhu:2023wci , one can further parametrize ${\bar{\nu}}$ in the form of
${\cal H}{\bar{\nu}}=\left[\alpha_{\bar{\nu}}(\tau)\left(\frac{k}{aM_{\rm LV}}\right)^{\beta_{\bar{\nu}}}\right]^{\prime},$ (2.8)
where $\beta_{\bar{\nu}}$ is an arbitrary number, $\alpha_{\bar{\nu}}$ is an arbitrary function of time, and $M_{\rm LV}$ denotes the energy scale of the Lorentz violation. The parameters $\alpha_{\bar{\nu}}$ and $\beta_{\bar{\nu}}$ depend on the specific modified theories of gravity.
In this paper, we consider the case with $\beta_{\bar{\nu}}=2$ and derive the corresponding modified waveform of GWs. We then explore the potential constraints on this modified waveform using proposed GW detectors such as aLIGO, CE, ET, Taiji, TianQin, and LISA. It is worth noting that this case can arise from the mixed temporal and spatial derivatives of the spacetime metric in the modified theories of gravity with spatial covariance Gao:2019liu ; Zhu:2022uoq ; Colombo:2014lta , for example, the mixed term in the form of
$\nabla_{k}K_{ij}\nabla^{k}K^{ij}$ (2.9)
can appear in both the spatial covariant gravity Gao:2019liu and Hořava-Lifshitz gravity Colombo:2014lta , where $\nabla_{k}$ denotes the covariant derivative associated with the spatial metric $g_{ij}$ and $K_{ij}$ is the extrinsic curvature tensor. It is also shown that by including mixed derivative terms, the non-protectable Hořava-Lifshitz gravity could be power-counting renormalizable and free of ghosts Colombo:2014lta . One can connect the coupling coefficient $c_{1}^{(2,2)}$ of Eq. (2.9) in modified gravities to the parameters $\alpha_{\bar{\nu}}$ and $M_{\rm LV}$ via
$\frac{\alpha_{\bar{\nu}}(\tau)}{M_{\rm LV}^{2}}=\frac{2c_{1}^{(2,2)}}{M_{\rm Pl}^{2}},$ (2.10)
where $M_{\rm Pl}$ is the reduced Planck mass. Considering the coupling coefficient as a constant, which is equivalent to treating $\alpha_{\bar{\nu}}$ as constant, we can simply set $\alpha_{\bar{\nu}}=1$ by absorbing its order of magnitude into $M_{\rm LV}$.
II.2 Amplitude modulation of GWs with Lorentz-violating damping effect
The nonzero parameter $\overline{\nu}$ provides a frequency-dependent damping of GW amplitudes during propagation. This means that GWs with different frequencies will experience different damping rates. This damping rate effect induces an amplitude modification in the GWs.
To study the modified waveform of GWs with this frequency-dependent damping of GW amplitudes, following the derivations in Refs. Zhao:2019xmm ; Qiao:2019wsh , let us decompose $h_{A}$ in Eq. (2.7) as
$h_{A}=h_{A}^{\rm GR}e^{-i\theta(\tau)},\;\;h_{A}^{\rm GR}={\cal A}_{A}^{\rm GR}e^{-i\Phi^{\rm GR}(\tau)}.$ (2.11)
Here $h_{A}^{\rm GR}$ is the solution of Eq. (2.7) when ${\bar{\nu}=0}$. ${\cal A}_{A}^{\rm GR}$ and $\Phi^{\rm GR}(\tau)$ are the amplitude and phase of $h_{A}^{\rm GR}$ respectively. With this decomposition, $\theta(\tau)$ encodes the correction arising from ${\bar{\nu}}$ which characterizes the Lorentz-violating damping effect.
Plugging the second equation of decomposition Eq. (2.11) into Eq. (2.7) with ${\bar{\nu}=0}$, one finds
$i\Phi^{\prime\prime}+\Phi^{\prime 2}+2i{\cal H}\Phi^{\prime}-k^{2}=0.$ (2.12)
Similarly, plugging the first equation of the decomposition Eq. (2.11) into Eq. (2.7), one gets
$i(\theta^{\prime\prime}+\Phi^{\prime\prime})+(\Phi^{\prime}+\theta^{\prime})^{2}+i(2+{\bar{\nu}}){\cal H}(\theta^{\prime}+\Phi^{\prime})-k^{2}=0.$
In GR, the time derivative of the phase $\Phi^{\prime}\sim k$. Here the wavenumber $k$ is connected to the frequency of GWs by $k=2\pi f/a_{0}$. Since the amplitude correction $\theta$ is induced by the expansion of the universe, one has $\theta^{\prime}\sim{\cal H}$ and $\theta^{\prime\prime}\sim{\cal H}^{2}$. Considering $k\gg{\cal H}$ and $\theta^{\prime\prime}\ll\Phi^{\prime}\theta^{\prime}\sim k\theta^{\prime}$, Eq. (II.2) can be simplified into
$2\theta^{\prime}+i{\cal H}{\bar{\nu}}\simeq 0$ (2.14)
Solving this equation gives
[…]
The Lorentz-violating damping effect in the phase $\theta$ is purely imaginary, indicating that it modifies the amplitude of the GWs during their propagation. Considering ${\bar{\nu}}$ is also frequency-dependent, such amplitude modulation depends on the frequency of GWs as well.
Specifically, with the above solution, one can write the waveform of GWs with Lorentz-violating effect as
$h_{A}=h_{A}^{\rm GR}e^{\delta h_{2}},$ (2.16)
where
[…]
$=$ $-\frac{1}{2}\left[\alpha_{\bar{\nu}}\left(\frac{k}{aM_{\rm LV}}\right)^{\beta_{\bar{\nu}}}\right]\Bigg{|}^{a_{0}}_{a_{e}}.$
It can be further rewritten in the form
$\delta h_{2}=-\frac{1}{2}\left(\frac{2u}{M_{\rm LV}{\cal M}}\right)^{\beta_{\bar{\nu}}}\Big{[}\alpha_{\bar{\nu}}(\tau_{0})-\alpha_{\bar{\nu}}(\tau_{e})(1+z)^{\beta_{\bar{\nu}}}\Big{]}.$
Here we define $u=\pi{\cal M}f$, where ${\cal M}=(1+z){\cal M}_{\rm c}$ represents the measured chirp mass, and ${\cal M}_{\rm c}\equiv(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ denotes the chirp mass of the binary system with component masses $m_{1}$ and $m_{2}$.
II.3 Amplitude modification to the waveform of GWs
To derive the modified waveform of GWs, we consider the GWs produced during the inspiral stage of the compact binaries. To directly contact with the observations, it is convenient to analyze the GWs in the Fourier domain. In this approach, under the stationary phase approximation, the responses of the detectors to the GW signal ${\tilde{h}}(f)$ can be written in the form of
${\tilde{h}}(f)=\Big{[}F_{+}h_{+}(f)+F_{\times}h_{\times}(f)\Big{]}e^{-2\pi if\Delta t},$ (2.19)
where $F_{+}$ and $F_{\times}$ denote the beam pattern functions of GW detectors, which depend on the GW source’s location and polarization angle Jaranowski:1998qm ; Zhao:2017cbb . The two polarizations of GWs, $h_{+}(f)$ and $h_{\times}(f)$, are related to the left- and right-handed polarization modes, $h_{\rm L}(f)$ and $h_{\rm R}(f)$, via
$h_{+}=\frac{h_{\rm L}+h_{\rm R}}{\sqrt{2}},\;\;h_{\times}=\frac{h_{\rm L}-h_{\rm R}}{\sqrt{2}i}.$ (2.20)
Then using Eq. (2.16), after tedious calculations, one obtains the following restricted form for the waveform of GWs in the Fourier domain as a function of the GW frequency $f$, i.e.,
$\tilde{h}(f)=\mathcal{A}_{\rm GR}f^{-7/6}e^{i\Psi_{\mathrm{GR}}(f)}e^{\delta h_{2}},$ (2.21)
where ${\cal A}_{\rm GR}$ and $\Psi_{\rm GR}$ represent the amplitude and phase of GWs of a compact binary inspiral signal in GR, and $e^{\delta h_{2}}$ with $\delta h_{2}$ being given by Eq. (II.2) is the amplitude correction to the waveform of GWs in GR. In the post-Newtonian approximation, the amplitude and the phase of GWs in GR can be expressed as Wang:2022yxb
${\cal A}_{\rm GR}=\frac{2}{5}\times\sqrt{\frac{5}{24}}\pi^{-2/3}\frac{\mathcal{M}^{5/6}}{D_{\rm L}},$ (2.22)
and
$\Psi_{\mathrm{GR}}(f)$ $=$ $2\pi ft_{\rm c}-\Phi_{\rm c}-\frac{\pi}{4}+\frac{3}{128\eta}u^{-5/3}\left\{1+\left(\frac{3715}{756}+\frac{55}{9}\eta\right)u^{2/3}-16\pi u\right.$ (2.23)
$+\left(\frac{15293365}{508032}+\frac{27145}{504}\eta+\frac{3085}{72}\eta^{2}\right)u^{4/3}+\pi\left(\frac{38645}{756}-\frac{65}{9}\eta\right)(1+\ln u)u^{5/3}$
$+\left[\frac{11583231236531}{4694215680}-\frac{640}{3}\pi^{2}-\frac{6848}{21}\gamma_{\mathrm{E}}-\frac{6848}{63}\ln(64u)+\left(-\frac{15737765635}{3048192}+\frac{2255}{12}\pi^{2}\right)\eta\right.$
$\left.\left.+\frac{76055}{1728}\eta^{2}-\frac{127825}{1296}\eta^{3}\right]u^{2}+\pi\left(\frac{77096675}{254016}+\frac{378515}{1512}\eta-\frac{74045}{756}\eta^{2}\right)u^{7/3}\right\},$
where the luminosity distance $D_{\rm L}$ is expressed as
$D_{\rm L}(z)=\frac{1+z}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{\mathrm{M}}(1+z^{\prime})^{3}+\Omega_{\Lambda}}}.$ (2.24)
Here we adopt the $\Lambda$CDM model with Hubble constant $H_{0}\approx 67.4$ km $\mathrm{s}^{-1}$ $\mathrm{Mpc}^{-1}$, matter density fractoin $\Omega_{\mathrm{m}}\approx 0.315$, and vacuum energy density fraction $\Omega_{\mathrm{\Lambda}}\approx 0.685$ Planck:2018vyg ; Zhang:2019ylr . $t_{c}$ and $\Phi_{c}$ are time and phase at coalescence, $\gamma_{\rm E}$ is the Euler constant, and $\eta=m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is the symmetric mass ratio.
Table 1: Characteristics of six GW detectors.
Detector Configuration $f_{\rm lower}$ [Hz] $f_{\rm upper}$ [Hz] Reference
aLIGO Rightangle 10 5000 Refs. Wang:2022yxb ; KAGRA:2023pio
ET Rightangle 1 10000 Ref. Hild:2010id
CE Rightangle 5 4000 Ref. Reitze:2019iox
Taiji Triangle 0.0001 0.1 Ref. Liu:2023qap
TianQin Triangle 0.0001 1 Ref. TianQin:2020hid
LISA Triangle 0.0001 0.1 Refs. Liu:2023qap ; LISA:2017pwj
III ANALYSIS FRAMEWORK WITH FISHER INFORMATION MATRIX
III.1 General considerations
In this section, we provide a brief overview of the matched-filter analysis with the FIM approach, which follows the method outlined for compact binary inspiral in Refs. Cutler:1994ys ; Finn:1992xs ; Poisson:1995ef . We calculate the noise-weighted inner product between the partial derivatives of each GW waveform parameter and the one-sided power spectral density (PSD) of the detector noises. This calculation yields the FIM. Inverting the FIM provides the variance-covariance matrix, where the diagonal elements represent the square root of the mean squared error for the estimated parameters of the signal. Previous studies have showcased the precision and utility of the FIM approach, particularly in situations with high signal-to-noise ratios (SNRs).
To be specific, we first give the noise-weighted inner product of two signals $h_{1}$ and $h_{2}$
[…]
where $\tilde{h}_{1}(f)$ and ${\tilde{h}}_{2}(f)$ are the Fourier transformation of GW signal $h(t)$, $S_{\rm n}(f)$ is the PSD of the detector’s noise, and the star superscript stands for complex conjugation. In the above expression, $f_{\rm max}\ (f_{\rm min})$ represents the instrumental maximum (minimum) threshold frequency.
For a given signal $h$, the SNR is defined as
$\rho=(h|h)^{1/2}.$ (3.2)
The modified waveform of GW from binary inspirals, influenced by the Lorentz-violating damping effect as described in Eq. (2.21), are generally characterized by a set of parameters $\theta_{i}$. In this context, one can define the FIM as
$F_{ij}=\left(\left.\frac{\partial\tilde{h}}{\partial\theta_{i}}\right|\frac{\partial\tilde{h}}{\partial\theta_{j}}\right).$ (3.3)
Here $\theta_{i}$ and $\theta_{j}$ represent the elements in the set of modified waveform parameters of GWs $\{\ln\mathcal{A},\ln\mathcal{M},\ln\mathcal{\eta},\phi_{\mathrm{c}},t_{\mathrm{c}},C_{\nu}\}$, where $C_{\nu}=M_{\mathrm{LV}}^{-2}$ characterizes the Lorentz-violating damping effect in the waveform. Then one can calculate each element of the FIM, which are given respectively in Appendix A.
In the large SNR approximation, if the noise is stationary and Gaussian, the probability that the GW signal $h(t)$ can be characterized by a given set of values of the parameters $\theta_{i}$,
$p(\theta_{i}|h)=p^{(0)}(\theta_{i})\exp\left[-\frac{1}{2}F_{ij}\Delta\theta^{i}\Delta\theta^{j}\right],$ (3.4)
where $p^{(0)}(\theta_{i})$ represents the distribution of prior information. Then, the standard deviations $\Delta\theta_{i}$ in measuring the parameter $\theta_{i}$, which mean $1\sigma$ bounds on parameters, can be calculated in the large SNR approximation. This can be obtained by taking the square root of the corresponding diagonal elements in the inverse of FIM,
$\Delta\theta_{i}=\sqrt{(F^{-1})_{ii}},$ (3.5)
where $F^{-1}$ is the inverse of FIM.
Our main purpose in employing FIM analysis here is to gauge the potential of future GW detectors in constraining the energy scale $M_{\rm{LV}}$ associated with Lorentz violation, which induces the Lorentz-violating damping effect in the propagation of GWs. We consider two types of GW detectors, ground-based GW detectors including aLIGO, CE, and ET, and space-based GW detectors, including Taiji, TianQin, and LISA. We summarize the information of all the six GW detectors in Table 1.
To investigate the variation tendency of $M_{\rm LV}$ in GW events, we choose the reasonable astrophysical horizon of each GW detector as the sources’ property and constrain $M_{\rm LV}$ by FIM. To maintain the effectiveness of FIM, the ranges of redshift $z$ and total mass $M$ are suitable enough to satisfy the SNR threshold ($\rho>8$ for ground-based detectors and $\rho>15$ for space-based detectors). For ground-based GW detectors, their detectable frequency bands are several Hz to thousand Hz. As the third generation GW detector, the frequency band of CE can reach $[5-4000]$ Hz. For space-based detectors, with huge arm-lengths so that they can detect the GWs from compact binary systems in a very low-frequency band $[10^{-4}-10^{-1}]$ Hz. The compact binary systems in this low-frequency band usually consist of supermassive black holes and their total mass range is $[10^{4}-10^{6}]\ \mathrm{M}_{\odot}$.
When calculating the integrals of the inner product, it is necessary to use the appropriate limits of integration. The minimum frequency $f_{\mathrm{min}}$ in Eq. (3.1) is taken as the instrumental minimum threshold frequency of the GW detector as shown in Table 1. The upper cut-off frequency $f_{\mathrm{max}}$ is chosen from $\mathrm{min}\{f_{\mathrm{upper}},f_{\mathrm{ISCO}}\}$, where $f_{\mathrm{upper}}$ is the upper frequency of detectors and $f_{\mathrm{ISCO}}$ is usually estimated by the innermost stable circular orbit (ISCO) Mirshekari:2011yq
$f_{\mathrm{ISCO}}=6^{-3/2}\pi^{-1}\eta^{3/5}\mathcal{M}^{-1}.$ (3.6)
When calculating the FIM, we set $C_{\nu}=0$ as the fiducial value. Additionally, the values of $t_{\rm c}$ and $\phi_{\rm c}$ do not impact the constraints on other parameters; hence, we set $t_{\rm c}=0$ and $\phi_{\rm c}=0$.
III.2 Noise power spectral density of detector
In Fig. 1, we illustrate the noise PSD of both ground-based and space-based detectors. The three ground-based detectors’ PSD are from the official data files quoted in Table 1. The other three PSD data are from the theoretical formula introduced below.
The sensitivity curve of Taiji can be obtained by the formula Liu:2023qap
$S_{\rm n}(f)$ $=\frac{10}{3L^{2}}\left(P_{\rm dp}+2(1+\mathrm{cos}^{2}(f/f_{*}))\frac{P_{\rm acc}}{(2\pi f)^{4}}\right)$
×$\times\left(1+0.6(\frac{f}{f_{*}})^{2}\right),$ (3.7)
where
$P_{\mathrm{dp}}=\left(8\times 10^{-12}\mathrm{~{}m}\right)^{2}\left(1+\left(\frac{2\,\mathrm{mHz}}{f}\right)^{4}\right)\mathrm{Hz}^{-1},$ (3.8)
$P_{\mathrm{acc}}$ $=\left(3\times 10^{-15}\mathrm{~{}m}\mathrm{~{}s}^{-2}\right)^{2}\left(1+\left(\frac{0.4\,\mathrm{mHz}}{f}\right)^{2}\right)$
×$\times\left(1+\left(\frac{f}{8\,\mathrm{mHz}}\right)^{4}\right)\mathrm{Hz}^{-1}.$ (3.9)
Here $P_{\rm dp}$ is the PSD of the displacement noise and $P_{\rm acc}$ is the PSD of the acceleration noise. $f_{*}=1/(2\pi L)$ and $L$ is the arm-length of the detector. For Taiji $L=3\times 10^{9}\ \rm m$.
For LISA, $L=2.5\times 10^{9}$ m and the displacement noise can be written as Liu:2023qap ; Berti:2004bd
$P_{\mathrm{dpL}}=\left(15\times 10^{-12}\mathrm{~{}m}\right)^{2}\left(1+\left(\frac{2\,\mathrm{mHz}}{f}\right)^{4}\right)\mathrm{Hz}^{-1}.$ (3.10)
The sensitivity curve for TianQin can be modeled by the following equation TianQin:2020hid
$S_{\rm n}(f)$ $=\frac{10}{3L^{2}}\left[S_{\rm x}+\frac{4S_{\rm a}}{(2\pi f)^{4}}\left(1+\frac{10^{-4}\mathrm{~{}Hz}}{f}\right)\right]$
×$\times\left[1+0.6\left(\frac{f}{f_{*}}\right)^{2}\right],$ (3.11)
where displacement measurement noise $S^{1/2}_{\rm x}$ and residual acceleration noise $S^{1/2}_{\rm a}$ are defined as
$S^{1/2}_{\rm x}=1\times 10^{-12}\ \mathrm{m}/\mathrm{Hz}^{1/2},$ (3.12)
and
$S^{1/2}_{\rm a}=1\times 10^{-15}\ \mathrm{m}\ \mathrm{s}^{-2}/\mathrm{Hz}^{1/2}.$ (3.13)
$f_{*}=c/(2\pi L)$ is the transfer frequency and $L=\sqrt{3}\times 10^{8}\ \mathrm{m}$.
IV RESULTS AND DISCUSSION
Table 2: The best constraints of the simulated single GW events and the combination of joint events from each GW detector. The results from the single event are the best constraint which are from Fig. 3. The redshift ranges and the total mass ranges are only for the joint events. The number of joint events contains 100 (for both ground-based and space-based detectors) simulated GW events. We choose $\rho>8$ and $\rho>15$ as the threshold for ground-based detectors and space-based detectors respectively.
Detector Single event [Gev] Joint redshift range Joint total mass range [$M_{\odot}$] Joint number Joint event [Gev]
aLIGO $8.54\times 10^{-22}$ [0.01-0.5] [3-10] 100 $2.39\times 10^{-21}$
CE $3.02\times 10^{-21}$ [0.01-0.5] [3-10] 100 $8.84\times 10^{-21}$
ET $2.39\times 10^{-21}$ [0.01-0.5] [3-10] 100 $7.18\times 10^{-21}$
Taiji $2.10\times 10^{-24}$ [5-10] $[4.2-9.2]\times 10^{3}$ 100 $5.86\times 10^{-24}$
TianQin $1.89\times 10^{-24}$ [0.01-5] $[1.5-2]\times 10^{4}$ 100 $4.56\times 10^{-24}$
LISA $1.52\times 10^{-24}$ [5-10] $[8.3-13.3]\times 10^{3}$ 100 $4.37\times 10^{-24}$
Figure 2: The results of $M_{\rm LV}$ and SNR of GW150914-like and GW170817-like events in three ground-based detectors. GW170817-like event is the compact binary system of two neutron stars and GW150914-like event is a system containing two black holes.
In this section, we present the results of the potential constraints on the parameters of the modified waveform of GWs with the Lorentz-violating damping effect. Among these parameters, we focus on the energy scale $M_{\rm LV}$ of Lorentz violation which characterizes the frequency-dependent damping of GWs during their propagations. To conduct a comprehensive and reliable analysis of GW parameter estimation, we consider the frequency bands of both ground-based and space-based detectors. We employ simulated GW data to constrain the Lorentz-violating damping effect through the FIM approach. For the simulated GW data, we refer to the astrophysical horizons of the detectors from Ref. Evans:2021gyd (for aLIGO, ET), Refs. Srivastava:2022slt ; Evans:2021gyd (for CE), Ref. Liu:2023qap (for Taiji), Ref. TianQin:2020hid (for TianQin), and Ref. LISA:2017pwj (for LISA). We set $\rho>8$ and $\rho>15$ as the thresholds for ground-based detectors and space-based detectors, respectively. To fully exploit the high event rates of future GW detectors, we also decide to conduct a multi-event joint constraint analysis. The results of constraining both individual GW events and their combinations are detailed in Table 2. The constraint results and SNR from GW150914-like and GW170817-like events are shown in Fig. 2. We depict the dependence of the lower bound of $M_{\mathrm{LV}}$ in Fig. 3.
Figure 3: Dependence of the lower bound of $M_{\rm LV}$ on the total mass and redshift of the binary inspiral systems for different detectors, aLIGO (top left), CE (top middle), ET (top and right), Taiji (bottom left), TianQin (bottom middle), and LISA (bottom right).
IV.1 $M_{\rm LV}$ from ground-based detectors
First, let us analyze the constraints on $M_{\rm LV}$ from the three ground-based detectors. In Fig. 2, we illustrate the constraints on $M_{\rm LV}$ (right panel) and SNRs (left panel) of two examples of GW events, the GW150914-like and GW170817-like events with three different ground-based detectors. As observed in the left panel of Fig. 2, the lower bounds on $M_{\rm LV}$ from both CE and ET extend to $\gtrsim 10^{-21}$ GeV. CE also gives the best constraints on $M_{\rm LV}$ for both GW150914-like and GW170817-like events. Notably, under similar redshift conditions, the GW170817-like event, characterized by a smaller total mass, attains a higher lower bound of $M_{\rm LV}$. Moving to the right panel of Fig. 2, CE exhibits a distinct advantage in SNR, suggesting its effectiveness in ensuring the detection of such GW events. There appears to be an indicative trend implying that GW events with a greater total mass could potentially yield higher SNRs.
In Table 2, it is evident that ground-based detectors, capable of observing high-frequency GWs, exhibit superior performance compared to their space-based counterparts. Among ground-based detectors, the strongest constraint for $M_{\rm LV}$ is achieved by the third-generation detector CE at $3.02\times 10^{-21}$ GeV. The result from the ET is marginally smaller than that of CE. For aLIGO, the constraint can reach $8.54\times 10^{-22}$ Gev.
Considering the prospect of observing a large number of GW events in the future, we also do research on the joint analysis for individual detectors. As anticipated, the outcomes from each detector exhibit a marked enhancement. The most notable improvement comes from CE, yielding the optimal result with $M_{\mathrm{LV}}>8.84\times 10^{-21}$ GeV. We combine 100 simulated GW events within the redshift range [0.01-0.5] Hz and total mass range [3-10] $M_{\odot}$ of sources. aLIGO attains a result of $M_{\mathrm{LV}}>2.39\times 10^{-21}\mathrm{GeV}$. This value is approximately twice as high as the result reported in Ref. Zhu:2023wci , which means that aLIGO continues to be a powerful tool in testing the Lorentz-violating damping effect. It is crucial to highlight that the arm-length of ET is 10 km. Meanwhile, CE is designed with arm-lengths of 20 km and 40 km. Due to their long arm-lengths and enhanced sensitivity, CE and ET can provide values of $M_{\mathrm{LV}}$ that are roughly more than double the value obtained by aLIGO. The expected detection rate, as reported in Refs. Li:2023gtu ; Han:2023exn , is conservatively estimated at 100.
IV.2 $M_{\rm LV}$ from space-based detectors
For the space-based detectors, including Taiji, TianQin, and LISA, the constraints on $M_{\rm LV}$ are about three orders of magnitude weaker than those from the ground-based detectors. As shown in Table 2, the constraints on $M_{\rm LV}$ is roughly at $\gtrsim 2\times 10^{-24}\ {\rm GeV}$ for single GW event. When considering a joint analysis of 100 simulated GW events, the constraints from all three detectors are roughly at $\gtrsim 5\times 10^{-24}\;{\rm GeV}$. Among the three detectors, Taiji achieves the best result, with $M_{\mathrm{LV}}>5.86\times 10^{-24}$ GeV. The reason why the constraints from the space-based detectors are weaker than those from the ground-based detectors is easy to understand. As one can see from Eq. (II.2), the amplitude correction to the waveform due to the Lorentz-violating damping effect is proportional to the square of the GW frequencies, which implies this effect is more sensitive to the higher GW frequencies, and thus the ground-based detectors can give stronger constraints than space-based detectors. We note that to estimate the number of events within our joint redshift and total mass range, we employ the data and methods from Refs. Wang:2021srv ; Klein:2015hvg .
IV.3 Trends of $M_{\rm LV}$
In Fig. 3, we illustrate how the lower bound of the Lorentz-violating parameter, $M_{\rm LV}$, varies with the total mass and redshift of binary inspiral systems across a selection of detectors. Specifically, for aLIGO, CE, and ET, we focus on a redshift range of $[0.01,0.5]$ Hz and a total mass range of $[3,50]\ M_{\odot}$. For Taiji, TianQin, and LISA, we extend the redshift range to $[0.01,10]$ Hz. The rationale behind selecting different total mass ranges for these latter three detectors is linked to ensuring that the SNR, $\rho$, exceeds 15 when the redshift value is maximized. This approach is designed to sharpen the visibility of the $M_{\mathrm{LV}}$ trend, highlighting the impact of high redshift values. As illustrated in Fig. 3, the highest values of the Lorentz violation energy scale, $M_{\mathrm{LV}}$, are achieved for sources that are both nearest and of the lowest mass, as observed in the cases of aLIGO, CE, ET, and TianQin. Conversely, for Taiji and LISA, the trend deviates. Here, $M_{\rm LV}$ does not attain its maximum in regions characterized by lower total masses and smaller redshifts. This discrepancy arises because the innermost stable circular orbit frequency, $f_{\mathrm{ISCO}}$, in this domain, surpasses the upper-frequency limit of these detectors. This is indicated by Eq. (3.6) and illustrated in Fig. 4. Consequently, this suggests that Taiji and LISA might be less efficient in detecting the final inspiral phase in compact binary systems that have lower redshifts and smaller total masses.
The choice to initiate the redshift ($z$) analysis from 0.01 stems from the observation that the lowest redshift value recorded in the LVK event catalog is 0.01. Considering that the mass of a neutron star is typically around 1.5 $M_{\odot}$, we adopt this figure as the minimum mass threshold for our total mass range. We cap the redshift at $z=0.5$ and set the maximum total mass at 50 $M_{\odot}$, which corresponds to the median GW source mass reported in the LVK events. To streamline our analysis, we concentrate on binary systems comprising either two black holes (BBH) or two neutron stars (BNS), assuming equal mass for both components.
V CONCLUSION
Figure 4: The distribution of $f_{\mathrm{ISCO}}$ for Taiji and LISA is depicted within the same astrophysical horizon, as illustrated in Fig. 3.
With the advent of future detectors, GWs are poised to play a pivotal role in testing gravity in the strong field regime. Both ground-based and space-based detectors are designed to capture GWs across different frequency bands, spanning from $10^{-4}$ to $10^{4}$ Hz. In this study, we delve into the investigation of the Lorentz-violating damping effect, which influences the propagation of GWs. We aim to evaluate the capability of both ground-based (aLIGO, CE, and ET) and space-based (Taiji, TianQin, and LISA) detectors in constraining this effect. We begin by formulating the modified equations of motion for the two polarizations of GWs. We then proceed to derive the altered GW waveform in the Fourier domain, incorporating the Lorentz-violating damping effect. Utilizing the FIM, we set out to quantify the constraints on the energy scale $M_{\mathrm{LV}}$, showcasing its projected sensitivity for each detector. For the FIM analysis, we establish detection thresholds for ground-based detectors and space-based detectors at $\rho>8$ and $\rho>15$, respectively. Additionally, we conduct a joint analysis of $M_{\mathrm{LV}}$ using simulated GW events to further our understanding of the constraints achievable with future GW observations.
For ground-based detectors, the tightest constraint from a single event is set by CE, with $M_{\rm LV}>3.02\times 10^{-21}$ GeV. When conducting a joint analysis of 100 GW events, this constraint improves to $M_{\rm LV}>8.84\times 10^{-21}$ GeV. Regarding space-based detectors, the results are in line with our expectations. For a single event, the constraints on $M_{\rm LV}$ from these detectors are approximately $\gtrsim 2\times 10^{-24}\;{\rm GeV}$, which is roughly three orders of magnitude less stringent than those obtained from ground-based detectors. Upon performing a joint analysis of 100 simulated GW events, the constraints from the space-based detectors converge to approximately $\gtrsim 5\times 10^{-24}\;{\rm GeV}$.
In our conservative estimation of event numbers, we note an improvement of more than twofold over the constraints obtained from individual events. We posit that, with the ongoing accumulation of observational data, even more stringent constraints on this effect will be achievable. Our analysis indicates that targeting the high-frequency band offers a more efficacious approach for constraining the Lorentz-violating damping effect. This suggests that ground-based detectors are more adept at imposing rigorous constraints on the effect in comparison to space-based detectors. Within this group of detectors, CE distinguishes itself by offering the most stringent lower bound on $M_{\mathrm{LV}}$.
Our analysis of the data distribution reveals that an effective strategy for aLIGO, CE, ET, and TianQin to constrain the Lorentz-violating damping effect involves concentrating on compact binary systems characterized by both a small total mass and low redshift. Conversely, for Taiji and LISA, targeting GW sources that have a small total mass but are situated at higher redshifts proves to be more appropriate. This strategic focus is informed by the differential sensitivity of these detectors to the frequency and amplitude of GW signals, which in turn affects their capability to place constraints on the Lorentz-violating effect.
ACKNOWLEDGMENTS
We thank Hong-Chao Zhang, Chao Zhang, and Peng-Ju Wu for useful discussions. Tao Zhu and Bo-Yang Zhang are supported by the National Key Research and Development Program of China under Grant No.2020YFC2201503, the National Natural Science Foundation of China under Grants No.12275238 and No. 11675143, the Zhejiang Provincial Natural Science Foundation of China under Grants No.LR21A050001 and No. LY20A050002, and the Fundamental Research Funds for the Provincial Universities of Zhejiang in China under Grant No. RF-A2019015. Jing-Fei Zhang and Xin Zhang are supported by the National Natural Science Foundation of China (Grants Nos. 11975072, 11875102, and 11835009), the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), and the National 111 Project (Grant No. B16009).
APPENDIX A: PARTIAL DERIVATIVES OF THE WAVEFORM OF BINARY INSPIRAL
In this appendix, we present the partial derivatives of the GW waveform parameters $\{\ln\mathcal{A},\ln\mathcal{M},\ln\mathcal{\eta},\phi_{\mathrm{c}},t_{\mathrm{c}},C_{\nu}\}$ as follows,
$\frac{\partial\tilde{h}(f)}{\partial\ln{\cal{A}}}$ $=$ $\tilde{h}(f),$ (A.1)
$\frac{\partial\tilde{h}(f)}{\partial\ln\cal{M}}$ $=$ $\frac{5}{6}\tilde{h}(f),$ (A.2)
$\frac{\partial\tilde{h}(f)}{\partial\ln{\eta}}$ $=$ $\frac{1}{2}+i\eta\Bigg{\{}-\frac{3}{128}\eta^{-2}u^{-5/3}C_{\Psi}+\frac{3}{128}\eta^{-1}u^{-5/3}\Bigg{[}\frac{55}{9}u^{2/3}+\left(\frac{27145}{504}+\frac{3085}{36}\eta\right)u^{4/3}-\frac{65\pi}{9}\left(1+\rm\ln{u}\right)u^{5/3}$ (A.3)
$+\left(-\frac{15737765635}{3048192}+\frac{2255}{12}\pi^{2}+\frac{76055}{864}\eta-\frac{127825}{432}\eta^{2}\right)u^{2}+\pi\left(\frac{378515}{1512}-\frac{74045}{378}\eta\right)u^{7/3}\Bigg{]}\Bigg{\}},$
$\frac{\partial\tilde{h}(f)}{\partial{\phi_{\rm c}}}$ $=$ $-i\tilde{h}(f),$ (A.4)
$\frac{\partial\tilde{h}(f)}{\partial{t_{\rm c}}}$ $=$ $(2\pi if)\tilde{h}(f),$ (A.5)
$\frac{\partial\tilde{h}(f)}{\partial{C_{\nu}}}$ $=$ $\Big{[}\frac{1}{2}(2\pi f)^{2}((1+z)^{2}-1)\Big{]}\tilde{h}(f),$ (A.6)
where $C_{\Psi}$ in Eq. (A.3) is given by
$C_{\Psi}$ $=$ $\left\{1+\left(\frac{3715}{756}+\frac{55}{9}\eta\right)u^{2/3}-16\pi u\right.$ (A.7)
$+\left(\frac{15293365}{508032}+\frac{27145}{504}\eta+\frac{3085}{72}\eta^{2}\right)u^{4/3}+\pi\left(\frac{38645}{756}-\frac{65}{9}\eta\right)(1+\ln u)u^{5/3}$
$+\left[\frac{11583231236531}{4694215680}-\frac{640}{3}\pi^{2}-\frac{6848}{21}\gamma_{\mathrm{E}}-\frac{6848}{63}\ln(64u)+\left(-\frac{15737765635}{3048192}+\frac{2255}{12}\pi^{2}\right)\eta\right.$
$\left.\left.+\frac{76055}{1728}\eta^{2}-\frac{127825}{1296}\eta^{3}\right]u^{2}+\pi\left(\frac{77096675}{254016}+\frac{378515}{1512}\eta-\frac{74045}{756}\eta^{2}\right)u^{7/3}\right\}.$